OpenStudy (anonymous):

Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius.

OpenStudy (anonymous):

chord*radius/2

OpenStudy (anonymous):

for area

OpenStudy (anonymous):

so its 8*8/2 =32

OpenStudy (anonymous):

sorry that's for sector

OpenStudy (anonymous):

you have to minus area of triangle from that

OpenStudy (anonymous):



OpenStudy (anonymous):

use A =(1/2) *base*height then Asegment =Asector -Atriangle

OpenStudy (anonymous):

Can you write out the formula for me?

OpenStudy (anonymous):

formula for>?

OpenStudy (anonymous):

Hold up I'm thinking.

OpenStudy (anonymous):

Draw the figure.

OpenStudy (anonymous):

you will get height = \[4\sqrt{3}\]

OpenStudy (anonymous):

 @RnR